Show that the simple pendulum is a special case of compound pendulum when the moment of inertia is mLsqure when the center of mass distance of compound pendulum equal to length of simple pendulum
1
Expert's answer
2020-06-30T18:23:56-0400

since, equation which compound pendulum is governed is

"\ddot{\theta}+\omega^2\theta=0"

Where,

"\omega^2=\frac{MgD}{I}"

Now, we have given "I=ML^2,D=L" ,thus "\omega^2=\frac{g}{L}" but equation of simple pendulum is given by

"\ddot{\theta}+\omega_1^2\theta=0"

where,

"\omega_1^2=\frac{g}{L}"

Thus,

"\omega=\omega_1"

Therefore, compound pendulum becomes simple pendulum.